The labor demand curve is L = 21 − w, where w denotes real wage. Wage negotiations are based on expected prices. At the beginning of a period trade unions and employers negotiate wages, which remain fixed for the entire period. Expected prices are P^e = 4 and nominal wages are W = 20. Determine the expected output for the (partial) production function Y (L) = 21L − 0.5L^2.
=> Y^e = ?
1
Expert's answer
2016-10-28T06:59:08-0400
L = 21 − w, P^e = 4, W = 20. Then the expected output for the (partial) production function Y (L) = 21L − 0.5L^2 = 21*(21 - 20) - 0.5*(21 - 20)^2 = 20.5 units, so P^e = 4 and Y^e = 20.5 units.